Circular No. 6137

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995E IN NGC 2441 P. Molaro, Trieste Observatory, reports the discovery by Alessandro

Gabrielcic, Agordo, Italy, of an supernova in NGC 2441 (α = 7^{h}52^{m}.2, δ = +73^{o}02′, equinox 2000.0), the new object being at mag V ∼ 15 on the CCD discovery frame taken on Feb. 20.85 UT. SN 1995E is located 7′′ east and 22′′ south of the galaxy’s center. Molaro, C. Boehm, and G. Tessicini further report that their spectrogram (range 500–900 nm, resolution FWHM ∼ 1.8 nm) taken with the 1.5-m Loiano Observatory telescope (+ BFOSC spectrograph) on Feb. 21.3 UT shows a strong absorption centered at 619 nm (FWHM 15.4 nm): “The presence of the absorption feature, which may be identified with Si II (635.5 nm), and the absence of pronounced Hα emission suggest that this is a type-Ia supernova. The expansion velocity amounts to −11000 km/s in the rest frame of NGC 2441, which is consistent with a type-Ia supernova close to maximum.”

SUPERNOVA 1995D IN NGC 2962 CCD magnitude from M. Blake, D. Lane, and P. Gray (Burke-Gaffney

Observatory, Saint Mary’s University, Halifax, NS): Feb. 14.149 UT, R = 13.7 ± 0.1 (comparison stars GSC 0242.0194 and 0242.0810). H. Mikuž, Ljubljana, Slovenia, reports the following photometry using glass filters (for coated CCDs; cf. Bessell 1990, PASP 102, 1181) that match the standard BVRI passbands: Feb. 19.89, V = 13.49 ± 0.02, B − V = +0.13 ± 0.03 (comparison star HR 3834). CCD magnitude estimate by S. Meister, Bulach, Switzerland: Feb. 20.99, 13.08 (comparison stars from Guide Star Catalogue).

Visual magnitude estimates: Feb. 17.833 UT, 12.8 (T. Vanmunster, Landen, Belgium); 18.903, 12.8 (P. Schmeer, Bischmisheim, Germany); 19.90, 13.4 (M. Villi and G. Cortini, Mt. Colombo, Italy); 21.9, 13.0 (L. Kiss, Szeged, Hungary); 22.083, 12.7 (Schmeer).

SU TAURI This R CrB-type variable has faded again, as indicated by the follow-

ing visual magnitude estimates: 1994 Aug. 14.123 UT, 11.7 (Schmeer); Dec. 15.104, 11.2 (Schmeer); 30.876, 11.5 (Schmeer); 1995 Jan. 5.806, 11.8 (Schmeer); 16.841, 12.3 (Schmeer); 30.918, 12.9 (Schmeer); Feb. 1.914, 13.7 (Schmeer); 19.8, 14.1 (L. Szentaskó, Veresegyház, Hungary).

1995 February 22 Daniel W. E. Green